Question 13 Not complete Marked out of 1.00 F Flag question metal Yellow light with a wavelength of 586.7 nm hits metal with a work function of 1.74 eV. What is the maximum kinetic energy of electrons emitted from this surface? (if no electrons are emitted type 0° as your answer.) eV Check Question 14 Not complete Marked out of 1 00 甲Fugquesion Which of the following changes would load to an increase in the maximum kinetic energy of the eijøcted electrons? Choose all of the correct options Select ono or more: Using red light instead of yellow light Increasing the intensity of the yellow light. Decreasing the work function of the metal Using blue light instead of yellow light. Check
Answer #1

13. The maximum kinetic energy of electrons emitted from the surface is, K_{max}=h\nu-\phi_o=\frac{hc}{\lambda}-\phi_o=\left(\frac{4.14\times 10^{-15}\times 3 \times 10^8}{586.7\times 10^{-9}}-1.74\right)\;eV\approx 0.38\; eV

14. Decreasing the work function of the metal

Using blue light instead of yellow light.